# Emmanuel Letellier Université Paris Cité, IMJ-PRG CNRS UMR 7586

### v2

@@ 2.2 Review on cohomology of $\mathrm{GL}_{n}$-character varieties
In [11] the authors defined a rational function $\mathbb{H}_{\bm{\mu}}(z,w)$ which conjecturally computes the polynomial $H_{c}(\mathcal{M}_{C};q,t)$. It is defined as follows.
Consider $k$ sets ${\bm{x}}_{1},\dots,{\bm{x}}_{k}$ of infinitely many independent variables and we let $\Lambda({\bm{x}}_{1},\dots,{\bm{x}}_{k})$ be the ring of functions separately symmetric in each of the set of variables. We then put
@@
Hence in the case $n=2$ Conjecture 2.2.2 reads.
@@
$H_{c}(\mathcal{M}_{C};q,t)=\frac{(q^{2}t^{4}+qt^{2})^{k}}{q^{2}t^{6}(q^{2}t^{2}-1)(q^{2}t^{4}-1)}+\frac{t^{2k-6}(q+1)^{k}}{(q^{2}-1)(q^{2}t^{2}-1)}-\frac{(2qt^{2})^{k-1}}{qt^{4}(q-1)(qt^{2}-1)}$
@@ 2.3 Twisted mixed Poincaré polynomials
Assume that $X$ is a $K$-variety endowed with a action of a finite group $W$. If $K=\overline{\mathbb{F}}_{q}$, we assume that the action of $W$ commutes with $F$. The group action preserves the weight filtration on $H_{c}^{\bullet}(X)$ and for $w\in W$, we define the $w$-twisted mixed Poincaré polynomial as
@@
The proof of this Theorem is very similar to that of [20, Theorem 2.8]. We give it below for completeness, see Proof 2.3.
@@
Given $(X/_{\overline{\mathbb{F}}_{q}},W)$ as above, for any $w\in W$, the map $wF:X\to X$ is a Frobenius morphism and gives thus another $\mathbb{F}_{q}$-structure on the $\overline{\mathbb{F}}_{q}$-variety $X$.
@@
If $K=\mathbb{C}$, we say that $(X/_{\mathbb{C}},W)$ has polynomial count with (twisted) counting polynomials $\{P_{w}(t)\}_{w\in W}$ if there exists a finitely generated $\mathbb{Z}$-subalgebra $R$ of $\mathbb{C}$, a separated finitely generated $R$-scheme $X_{R}$ equipped with a $W$-action which gives back $X$ with its $W$-action after scalar extensions from $R$ to $\mathbb{C}$, such that for any ring homomorphism $\varphi:R\to\mathbb{F}_{q}$, $(X^{\varphi},W)$ has polynomial counting with (twisted) counting polynomials $\{P_{w}(t)\}_{w\in W}$.
@@
Theorem 2.3.3 above can be deduced from Theorem 2.3.1 as follows. Consider a variety $X/_{\mathbb{C}}$ and $R$ as above and $U\subseteq\operatorname{Spec}(R)$ as in section 2.1.1. Since the action of $W$ is defined over $R$, the isomorphism (2.1.3) is compatible with the $W$-action.
@@
If $P_{w}(t)=\sum_{i}t^{i}c_{w,i}$, from eq.(2.3.3) we deduce that we have
@@
From [20, Lemma 2.9], we deduce that we have
@@
From Formula 2.1.4, it is enough to show the case of $K=\overline{\mathbb{F}}_{q}$. Take thus a finite field $\mathbb{F}_{q}$ of characteristic $\neq 2$ and consider a generic $k$-tuple $C=(C_{1},\dots,C_{k})$ of semisimple regular conjugacy classes of $\operatorname{GL}_{2}(\overline{\mathbb{F}}_{q})$ with eigenvalues in $\mathbb{F}_{q}$ and such that the first $m$ conjugacy classes are the conjugacy class with eigenvalues $1,-1$.
Then group $H_{m}$ acts on $\mathcal{M}_{C}$ by multiplication coordinate by coordinate on the first $m$-th coordinates.
@@
Fix $i=1,\dots,m$, let $\alpha_{i}\in\overline{\mathbb{F}}_{q}^{*}$ be defined as $\alpha_{i}=1$ if $y_{i}=1$ and as an element of $\mathbb{F}_{q^{2}}^{*}$ such that $\alpha_{i}^{q}=-\alpha_{i}$.
@@
Notice that $\alpha_{1}\cdots\alpha_{m}\in\mathbb{F}_{q}^{\times}$ as there is an even number of non-trivial $y_{i}$.
@@
Now for an $F$-stable conjugacy class ${O}$ of $\operatorname{GL}_{2}(\overline{\mathbb{F}}_{q})$ denote by $1_{\bm{O}}:\operatorname{GL}_{2}(\mathbb{F}_{q})\rightarrow\mathbb{C}^{*}$ the function that takes the value $1$ at ${\bm{O}}^{F}$ and 00 elsewhere.
@@
It follows from [18, Theorem 4.14] that the RHS equals to
@@
Indeed, since the conjugacy classes ${O}_{i}$ are semisimple, the function ${\bm{X}}_{\overline{O}_{i}}$ in [18, Theorem 4.14] is our function $1_{{\bm{O}}_{i}}$. In loc. cit. we introduced the notion of type of a conjugacy class of $\operatorname{GL}_{n}(\mathbb{F}_{q})$ (see [18, §4.1.2]) which we could avoid here as we are working with semisimple conjugacy classes of $\operatorname{GL}_{2}$. The type of a split semisimple regular conjugacy class of $\operatorname{GL}_{2}(\mathbb{F}_{q})$ is $\omega=(1,1)(1,1)$ and the symmetric function $s_{\omega}$ (see [18, Remark 4.1]) is our symmetric function $h_{(1^{2})}$. The type of a non-split semisimple regular conjugacy class of $\operatorname{GL}_{2}(\mathbb{F}_{q})$ like our conjugacy class ${O}_{i}$ with $y_{i}=-1$ (it has eigenvalues in $\mathbb{F}_{q^{2}}-\mathbb{F}_{q}$) is $\omega=(2,1)$ and $s_{\omega}$ is our symmetric function $p_{2}$. Finally in the formula for $\mathbb{H}_{\bm{\omega}}(z,w)$ in [18, Above Lemma 4.2], there is a sign which in our case is always $1$ as we have an even number of twisted conjugacy classes. ∎
3 $\mathrm{PGL}_{2}$-character varieties
3.1 Local systems on semisimple conjugacy classes of $\mathrm{PGL}_{2}$
@@
has two connected components and so there are exactly two (up to isomorphism) irreducible $\operatorname{PGL}_{2}(K)$-equivariant local systems on the $\operatorname{PGL}_{2}(K)$-conjugacy class $\mathcal{C}_{-1}$ of $p(g_{-1})$. The non-trivial local system $\mathcal{L}$ is obtained from the $2$-covering
@@
$p_{*}(\kappa)\cong\kappa\oplus\mathcal{L}.$
@@
The local system $\mathcal{L}$ has thus the following explicit description. For an open $U\subseteq\mathcal{C}_{-1}$, we have
$\mathcal{L}(U)=\{f\in\kappa(p^{-1}(U))\ |\ f\circ\sigma=-f\}.$ (3.1.1)
@@
3.2 Geometry of $\mathrm{PGL}_{2}$-character stacks and varieties
Let $p^{k}:\operatorname{GL}_{2}^{k}\to\operatorname{PGL}_{n}^{k}$ be the $k$-th product of the projection map. Fix a $k$-tuple $\mathcal{C}=(\mathcal{C}_{1},\dots,\mathcal{C}_{k})$ of regular semisimple conjugacy classes of $\operatorname{PGL}_{2}$ and consider a $k$-tuple $C=(C_{1},\dots,C_{k})$ of conjugacy classes of $\operatorname{GL}_{2}$ such that $p(C_{i})=\mathcal{C}_{i}$ for each $i=1,\dots,k$. Fix a square root
@@ 3.2 Geometry of $\mathrm{PGL}_{2}$-character stacks and varieties
In [13, Definition 9], the authors give a definition of a generic $k$-tuple of conjugacy classes for any reductive group $G$. It is not hard to see that their definition agrees with our definition in the case of $\operatorname{PGL}_{2}$.
@@
If $\mathcal{C}$ is generic, the $\operatorname{PGL}_{2}$-character stack $\mathcal{M}_{\mathcal{C}}$ is a smooth algebraic variety (i.e. the canonical morphism $\mathcal{M}_{\mathcal{C}}\rightarrow M_{\mathcal{C}}$ is an isomorphism of stacks) of dimension $62k-$ 6.
@@ 3.3 Local systems on $\mathrm{PGL}_{2}$-character stacks
$\mathcal{E}_{i}=\begin{cases}\mathcal{L}&\text{ if }i\in\{1\}\cup B,\\ \kappa&\text{ otherwise}.\end{cases}$
@@
$\mathcal{L}_{B}=(i_{\mathcal{C}})^{*}(\mathcal{L}\boxtimes\kappa\boxtimes\cdots\boxtimes\kappa)$
@@
3.4 Cohomology of generic $\mathrm{PGL}_{2}$-character varieties
@@
Notice that, for such a $\mathcal{C}$, we can find a $k$-tuple $C$ of conjugacy classes of $\operatorname{GL}_{2}(\overline{\mathbb{F}}_{q})$ which are $F$-stable and split semisimple and such that $p(C_{i})=\mathcal{C}_{i}$. For instance we can take $C_{i}$ to be the conjugacy class of $g_{x_{i}}$. In this case, we assume in what follows that $\lambda_{C}\in\mathbb{F}_{q}^{*}$, or, equivalently, that
@@ 4.1 Deligne-Lusztig theory for $\mathrm{GL}_{2}$ and $\mathrm{SL}_{2}$
The affine algebraic variety $L^{-1}(U)$ is equipped with an action of $G^{F}$ by left multiplication and with an action of $T^{F}$ by right multiplication (because $U$ is a normal subgroup of $B$). These actions induce actions on the compactly supported $\ell$-adic cohomology groups $H_{c}^{i}(L^{-1}(U),\overline{{\mathbb{Q}}})$.
@@
When $G=\operatorname{GL}_{n}$, we can can express all irreducible characters of $G^{F}$ as a linear combination of these Deligne-Lusztig characters (this is not the case for $\operatorname{SL}_{2}$ as we will see below).
When the maximal torus $T$ is contains in some $F$-stable Borel subgroup $B$ (we say that $T$ is split), then $R_{T}^{G}(\theta)$ as the following more concrete description (see [9, below Definition 11.1]):
@@
[9, Proposition 13.22] Let $T$ be an $F$-stable maximal torus of $\operatorname{GL}_{n}$ and denote by $T^{\prime}$ the maximal torus $T\cap\operatorname{SL}_{n}$ of $\operatorname{SL}_{n}$. Let $\theta$ be a linear character of $T^{F}$, then
@@
The following table gives the values of Deligne-Lusztig characters $R_{T_{w}}^{G}(\theta)$ in the case of $G=\operatorname{GL}_{2}$ (and by restriction we get also the values in the $G=\operatorname{SL}_{2}$ case), see [9, §15.9]
@@
The character tables of $\operatorname{GL}_{2}(\mathbb{F}_{q})$ and $\operatorname{SL}_{2}(\mathbb{F}_{q})$ from Deligne-Lusztig theory can be found for instance in [9, 15.9].

### v3

@@ 2.2 Review on cohomology of $\mathrm{GL}_{n}$-character varieties
In [11, Section 2.3] the authors defined a rational function $\mathbb{H}_{\bm{\mu}}(z,w)$ which conjecturally computes the polynomial $H_{c}(\mathcal{M}_{C};q,t)$. It is defined as follows.
Consider $k$ sets ${\bm{x}}_{1},\dots,{\bm{x}}_{k}$ of infinitely many independent variables and let $\Lambda({\bm{x}}_{1},\dots,{\bm{x}}_{k})$ be the ring of functions separately symmetric in each of the set of variables. We then put
@@
Hence in the case $n=2$, Conjecture 2.2.2 reads.
@@
$H_{c}(\mathcal{M}_{C};q,t)=\frac{(q^{2}t^{4}+qt^{2})^{k}}{q^{2}t^{6}(q^{2}t^{2}-1)(q^{2}t^{4}-1)}+\frac{t^{2k-6}(q+1)^{k}}{(q^{2}-1)(q^{2}t^{2}-1)}-\frac{(2qt^{2})^{k-1}}{qt^{4}(q-1)(qt^{2}-1)}.$
@@ 2.3 Twisted mixed Poincaré polynomials
Assume that $X$ is a $K$-variety endowed with an action of a finite group $W$. If $K=\overline{\mathbb{F}}_{q}$, we assume that the action of $W$ commutes with $F$. The group action preserves the weight filtration on $H_{c}^{\bullet}(X)$ and for $w\in W$, we define the $w$-twisted mixed Poincaré polynomial as
@@
The proof of this Theorem is very similar to that of [19, Theorem 2.8]. We give it below after Theorem 2.3.3.
@@
Given $(X/_{\overline{\mathbb{F}}_{q}},W)$ as above, for any $w\in W$, the map $wF:X\to X$ is a Frobenius morphism and gives thus another $\mathbb{F}_{q}$-structure of the $\overline{\mathbb{F}}_{q}$-variety $X$.
@@
If $K=\mathbb{C}$, we say that $(X/_{\mathbb{C}},W)$ has polynomial count with (twisted) counting polynomials $\{P_{w}(t)\}_{w\in W}$ if there exists a finitely generated $\mathbb{Z}$-subalgebra $R$ of $\mathbb{C}$, a separated $R$-scheme $X_{R}$ equipped with a $W$-action which gives back $X$ with its $W$-action after scalar extensions from $R$ to $\mathbb{C}$, such that for any ring homomorphism $\varphi:R\to\mathbb{F}_{q}$, $(X^{\varphi},W)$ has polynomial counting with (twisted) counting polynomials $\{P_{w}(t)\}_{w\in W}$.
@@
Theorem 2.3.3 above can be deduced from Theorem 2.3.1 as follows. Consider a variety $X/_{\mathbb{C}}$ and $R$ as above and $U\subseteq\operatorname{Spec}(R)$ as in section 2.1. Since the action of $W$ is defined over $R$, the isomorphism (2.1.3) is compatible with the $W$-action.
@@
If $P_{w}(t)=\sum_{i}t^{i}c_{w,i}$, from Formula (2.3.3) we deduce that
@@
From [19, Lemma 2.9], we deduce that we have
@@
From Formula 2.1.4, it is enough to show the case of $K=\overline{\mathbb{F}}_{q}$. Take thus a finite field $\mathbb{F}_{q}$ of characteristic $\neq 2$ and consider a generic $k$-tuple $C=(C_{1},\dots,C_{k})$ of semisimple regular conjugacy classes of $\operatorname{GL}_{2}(\overline{\mathbb{F}}_{q})$ with eigenvalues in $\mathbb{F}_{q}^{*}$ and such that the first $m$ conjugacy classes are the conjugacy class with eigenvalues $1,-1$.
Then the group $H_{m}$ acts on $\mathcal{M}_{C}$ by multiplication coordinate by coordinate on the first $m$-th coordinates.
@@
Fix $i=1,\dots,m$, let $\alpha_{i}\in\overline{\mathbb{F}}_{q}^{*}$ be defined as $\alpha_{i}=1$ if $y_{i}=1$ and as an element of $\mathbb{F}_{q^{2}}^{*}$ such that $\alpha_{i}^{q}=-\alpha_{i}$ if $y_{i}=-1$.
@@
Notice that $\alpha_{1}\cdots\alpha_{m}\in\mathbb{F}_{q}^{*}$ as there is an even number of non-trivial $y_{i}$.
@@
Now for an $F$-stable conjugacy class ${O}$ of $\operatorname{GL}_{2}(\overline{\mathbb{F}}_{q})$ denote by $1_{O}:\operatorname{GL}_{2}(\mathbb{F}_{q})\rightarrow\mathbb{C}^{*}$ the function that takes the value $1$ at ${O}^{F}$ and 00 elsewhere.
@@
It follows from [17, Theorem 4.14] that the RHS equals to
@@
Indeed, since the conjugacy classes ${O}_{i}$ are semisimple, the function ${\bm{X}}_{\overline{O}_{i}}$ in [17, Theorem 4.14] is our function $1_{{O}_{i}}$. In loc. cit. we introduced the notion of type of a conjugacy class of $\operatorname{GL}_{n}(\mathbb{F}_{q})$ (see [17, §4.1.2]) which we could avoid here as we are working with semisimple conjugacy classes of $\operatorname{GL}_{2}$. The type of a split semisimple regular conjugacy class of $\operatorname{GL}_{2}(\mathbb{F}_{q})$ is $\omega=(1,1)(1,1)$ and the symmetric function $s_{\omega}$ (see [17, Remark 4.1]) is our symmetric function $h_{(1^{2})}$. The type of a non-split semisimple regular conjugacy class of $\operatorname{GL}_{2}(\mathbb{F}_{q})$ like our conjugacy class ${O}_{i}$ with $y_{i}=-1$ (it has eigenvalues in $\mathbb{F}_{q^{2}}-\mathbb{F}_{q}$) is $\omega=(2,1)$ and $s_{\omega}$ is our symmetric function $p_{2}$. Finally in the formula for $\mathbb{H}_{\bm{\omega}}(z,w)$ in [17, Above Lemma 4.2], there is a sign which in our case is always $1$ as we have an even number of twisted conjugacy classes. ∎
3 $\mathrm{PGL}_{2}$-character stacks
@@
3.1 Local systems on semisimple conjugacy classes of $\mathrm{PGL}_{2}(K)$
@@
has two connected components and so there are exactly two (up to isomorphism) irreducible $\operatorname{PGL}_{2}(K)$-equivariant local systems on the $\operatorname{PGL}_{2}(K)$-conjugacy class $\mathcal{C}_{-1}$ of $p(g_{-1})$. The non-trivial local system $\mathcal{L}_{\epsilon}$ is obtained from the $2$-covering
@@
$p_{*}(\kappa)\cong\kappa\oplus\mathcal{L}_{\epsilon}.$
@@
The local system $\mathcal{L}_{\epsilon}$ has thus the following explicit description. For an open $U\subseteq\mathcal{C}_{-1}$, we have
$\mathcal{L}_{\epsilon}(U)=\{f\in\kappa(p^{-1}(U))\ |\ f\circ\sigma=-f\}.$ (3.1.1)
@@
3.2 Geometry of $\mathrm{PGL}_{2}$-character stacks
Let $p^{k}:(\operatorname{GL}_{2})^{k}\to(\operatorname{PGL}_{2})^{k}$ where $p:\operatorname{GL}_{2}\rightarrow\operatorname{PGL}_{2}$ is the quotient map. Fix a $k$-tuple $\mathcal{C}=(\mathcal{C}_{1},\dots,\mathcal{C}_{k})$ of regular semisimple conjugacy classes of $\operatorname{PGL}_{2}(K)$ and consider a $k$-tuple $C=(C_{1},\dots,C_{k})$ of conjugacy classes of $\operatorname{GL}_{2}(K)$ such that $p(C_{i})=\mathcal{C}_{i}$ for each $i=1,\dots,k$. Fix a square root
@@ 3.2 Geometry of $\mathrm{PGL}_{2}$-character stacks
In [9, Definition 9], the authors give a definition of a generic $k$-tuple of conjugacy classes for any reductive group $G$. It is not hard to see that their definition agrees with our definition in the case of $\operatorname{PGL}_{2}$.
@@
(1) If $\mathcal{C}$ is generic and $\mathcal{C}_{1},\dots,\mathcal{C}_{k}$ are non degenerate, the $\operatorname{PGL}_{2}$-character stack $\mathcal{M}_{\mathcal{C}}$ is a smooth algebraic variety (i.e. the canonical morphism $\mathcal{M}_{\mathcal{C}}\rightarrow M_{\mathcal{C}}$ is an isomorphism of stacks) of dimension $62k-$ 6.
(2) If $\mathcal{C}$ is generic and there are $m$ degenerate classes among $\mathcal{C}_{1},\dots,\mathcal{C}_{k}$ with $m\geq 1$, the $\operatorname{PGL}_{2}$-character stack $\mathcal{M}_{\mathcal{C}}$ is a smooth Deligne-Mumford stack of dimension $62k-$ 6. More precisely, we have an isomorphism $\mathcal{M}_{\mathcal{C}}\cong[\mathcal{M}_{C}^{+}/H_{m}]$.
Remark 3.2.7.
If there are $m$ degenerate classes between $\mathcal{C}_{1},\dots,\mathcal{C}_{k}$ with $m\geq 1$, it is not always true that the $\operatorname{PGL}_{2}$-action on $X_{\mathcal{C}}$ is free, i.e. that the canonical morphism $\mathcal{M}_{\mathcal{C}}\to M_{\mathcal{C}}$ is an isomorphism. Consider for example the case of $k=5$, $m=4$ and where $\mathcal{C}_{5}$ is the $\operatorname{PGL}_{2}$-conjugacy class of $p(g_{4})$. Pick $\lambda_{\mathcal{C}}=2$.
It is not hard to see that the $5$-tuple $\mathcal{C}$ is generic. We have that
$(X_{1},\dots,X_{5})=\left(\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},\begin{pmatrix}0&\frac{1}{2}\\ 2&0\end{pmatrix},\begin{pmatrix}-1&0\\ 0&1\end{pmatrix},\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\begin{pmatrix}4&0\\ 0&1\end{pmatrix}\right)\in X^{+}_{C}$
and therefore $(p(X_{1}),\dots,p(X_{5}))\in X_{\mathcal{C}}$. Notice that
$g_{-1}\cdot(X_{1},\dots,X_{5})=(X_{1},-X_{2},X_{3},-X_{4},X_{5})$ (3.2.7)
from which we deduce that $p(g_{-1})\cdot(p(X_{1}),\dots,p(X_{5}))=(p(X_{1}),\dots,p(X_{5}))$, i.e. $p(g_{-1})\in\operatorname{Stab}_{\operatorname{PGL}_{2}}((p(X_{1}),\dots,p(X_{5})))$.
@@ 3.3 Local systems on $\mathrm{PGL}_{2}$-character stacks
$\mathcal{E}_{i}=\begin{cases}\mathcal{L}_{\epsilon}&\text{ if }i\in\{1\}\cup B,\\ \kappa&\text{ otherwise}.\end{cases}$
@@
$\mathcal{L}_{B}=(i_{\mathcal{C}})^{*}(\mathcal{L}_{\epsilon}\boxtimes\kappa\boxtimes\cdots\boxtimes\kappa)$
@@
3.4 Cohomology of generic $\mathrm{PGL}_{2}$-character stacks
@@
Notice that, for such a $\mathcal{C}$, we can find a $k$-tuple $C$ of semisimple conjugacy classes of $\operatorname{GL}_{2}(\overline{\mathbb{F}}_{q})$ which are $F$-stable and split and such that $p(C_{i})=\mathcal{C}_{i}$. For instance we can take $C_{i}$ to be the conjugacy class of $g_{x_{i}}$ for each $i$. In this case, we assume in what follows that $\lambda_{C}\in\mathbb{F}_{q}^{*}$, or, equivalently, that
@@ 3.4.1 The non-degenerate case
Remark 3.4.3.
We know from Theorem 2.2.3 and Theorem 2.2.5 that the coefficient of the highest power of $q$ in $\mathbb{H}_{\bm{\mu}}\left(\frac{1}{\sqrt{q}},\sqrt{q}\right)$ equals $1$. Therefore the coefficient of the highest power of $q$ in $E(\mathcal{M}_{\mathcal{C}};q)$ equals $2$ which is also the number of connected components of the center of the dual group $\operatorname{SL}_{2}$ of $\operatorname{PGL}_{2}$. This has been previously observed for an arbitrary connected reductive group [9, Remark 3 (iii)].
@@ 4.1 Deligne-Lusztig theory for $\mathrm{GL}_{2}$ and $\mathrm{SL}_{2}$
The affine algebraic variety $L^{-1}(U)$ is equipped with an action of $G^{F}$ by left multiplication and with an action of $T^{F}$ by right multiplication (because $U$ is a normal subgroup of $B$). These actions induce actions on the compactly supported $\ell$-adic cohomology groups $H_{c}^{i}(L^{-1}(U),\overline{{\mathbb{Q}}}_{\ell})$.
@@
When $G=\operatorname{GL}_{n}$, we can express all irreducible characters of $G^{F}$ as a linear combination of these Deligne-Lusztig characters (this is not the case for $\operatorname{SL}_{2}$ as we will see below).
When the maximal torus $T$ is contained in some $F$-stable Borel subgroup $B$ (we say that $T$ is split), then $R_{T}^{G}(\theta)$ has the following more concrete description (see [8, below Definition 11.1]):
@@
[8, Proposition 13.22] Let $T$ be an $F$-stable maximal torus of $\operatorname{GL}_{n}$ and denote by $T^{\prime}$ the maximal torus $T\cap\operatorname{SL}_{n}$ of $\operatorname{SL}_{n}$. Let $\theta$ be a linear character of $T^{F}$, then
@@
The following table gives the values of Deligne-Lusztig characters $R_{T_{w}}^{G}(\theta)$ in the case of $G=\operatorname{GL}_{2}$ (and by restriction we get also the values in the $G=\operatorname{SL}_{2}$ case), see [8, §15.9]
@@
The character tables of $\operatorname{GL}_{2}(\mathbb{F}_{q})$ and $\operatorname{SL}_{2}(\mathbb{F}_{q})$ from Deligne-Lusztig theory can be found for instance in [8, 15.9].